OpenStudy (anonymous):

Let logP/N=5 and logM/N=9 what is the relationship between P and M

myininaya (myininaya):

\[\log(\frac{P}{N})=5 ; \log(\frac{M}{N})=9\] Did you try expanding both left hand sides of the equations

myininaya (myininaya):

\[\log(P)-\log(N)=5 ; \log(M)-\log(N)=9\]

myininaya (myininaya):

solve both ln(N)

myininaya (myininaya):

for ln(N)

myininaya (myininaya):

or log(N) whatever you want to call it

myininaya (myininaya):

then set the equations equal

myininaya (myininaya):

\[\log(P)=5+\log(N) ; \log(M)=9+\log(N)\] \[\log(P)-5=\log(N) ; \log(M)-9=\log(N)\]

myininaya (myininaya):

so what does this imply?

OpenStudy (anonymous):

A.m=p4 b. m=10000p c. m=0.0001p d. p=10000 those are the answers given...

myininaya (myininaya):

\[\log(P)-5=\log(M)-9\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

it was b